OpenStudy (anonymous):

Prove 9n - 1 is divisible by 8. The first two steps will be provided. For step 3, match the right and left sides of the equation to prove the statement. Step 1: When n = 1, 9n -1 = 91 - 1 = 8. Eight is divisible by 8, so the equation is true for n = 1. Step 2: Assume that 9k-1 is divisible by 8 for some positive integer, k. This means that there is a whole number r such that 9k-1 = 8r. Step 3: Show that the statement n = k + 1 is true. 9k - 1 = 9k = 9(9k) = 9k+1 = 9k+1 - 1 = 9k+1 - 1 = @mathmale @YanaSidlinskiy I am so confused!

OpenStudy (midhun.madhu1987):

I'll attach a file that I have written to prove another question.. You can compare it with this.. let me know if you need any assistance...

OpenStudy (anonymous):

You have to,now,prove, what @YanaSidlinskiy will tell you.. :P

OpenStudy (helder_edwin):

R u sure about the statement? because it is false.

OpenStudy (helder_edwin):

say u set n=2 then \[\large 9n-1=9\cdot2-1=18-1=17 \] and 17 is not divisble by 8.